OpenStudy (anonymous):

find the instantaneous rate of change of w with respect to z if w=7/3z^2. Show step by step please

OpenStudy (anonymous):

Take the derivative with respect to x: \[\frac{dw}{dz}=\frac{7}{3} \frac{d}{dz}z^{-2}; \frac{d}{dz}z^n=n * z^{n-1}; \implies \frac{7}{3} \frac{d}{dz} z^{-2}=\frac{-14}{3}z^{-3}=\frac{-14}{3 z^3}\]

OpenStudy (anonymous):

thank you